OpenStudy (friend736):

For the two - variable equation y - 3x = 15 find the matching y value when x = 1/3

OpenStudy (friend736):

Please help i give medals

Vocaloid (vocaloid):

x = 1/3 means that we can replace "x" with "1/3" whenever x shows up

Vocaloid (vocaloid):

so, if y - 3x = 15 and x = 1/3 what is y equal to?

OpenStudy (friend736):

i dont know im having trouble:(

Vocaloid (vocaloid):

easier question if x = 1/3, what is 3x?

OpenStudy (skullpatrol):

The statement: $$ x= \dfrac{1}{3}$$ means that wherever you see the variable x you can replace it with the number 1/3. This is called the substitution principle: An expression maybe replaced by another expression that has the same value. So, x = 1/3 is saying that x and 1/3 have the same value, that means x can be replaced by 1/3.

OpenStudy (skullpatrol):

This is why we can replace 4 + 2 with 6 in the statement 4 + 2 = 6 and get the true statement 6 = 6, because 4 + 2 and 6 have the same value.